# GPT-Fathom: Benchmarking Large Language Models to Decipher the Evolutionary Path towards GPT-4 and Beyond

### v5

@@ 1 Introduction
  • Inconsistent settings: The evaluation settings, such as the number of in-context example “shots”, whether Chain-of-Thought (CoT; Wei et al. 2022) prompting is used, methods of answer parsing and metric computation, etc., often differ across the existing LLM works. Moreover, most of the released LLMs do not disclose their prompts used for evaluation, making it difficult to reproduce the reported scores. Different settings and prompts may lead to very different evaluation results, which may easily skew the observations. Yet, many existing LLM leaderboards reference scores from other papers without consistent settings and prompts, which may inadvertently encourage cherry-picking favored settings and prompts for better results. To achieve reliable conclusions, it is crucial to make apples-to-apples LLM comparisons with consistent settings and prompts.
@@
Benchmarks constantly play a pivotal role in steering the evolution of AI and, of course, directing the advancement of LLMs as well. There are many great existing LLM evaluation suites. By comparing GPT-Fathom with previous works, we summarize the major difference as follows: 1) HELM (Liang et al., 2023) primarily uses answer-only prompting (without CoT) and has not included the latest leading models such as GPT-4 (as of the time of writing); 2) Open LLM Leaderboard (Beeching et al., 2023) focuses on open-source LLMs, while we jointly consider leading closed-source and open-source LLMs; 3) OpenCompass (Contributors, 2023) evaluates latest open-source and closed-source LLMs (all released after 2023/03), while we cover both leading LLMs and OpenAI’s earlier models to decipher the evolutionary path from GPT-3 to GPT-4; 4) InstructEval (Chia et al., 2023) is designed for evaluating instruction-tuned LLMs, while we evaluate both base and SFT / RLHF models; 5) AlpacaEval (Li et al., 2023) evaluates on simple instruction-following tasks as a quick and cheap proxy of human evaluation, while we provide systematic evaluation of various aspects of LLM capabilities; 6) Chatbot Arena (Zheng et al., 2023) evaluates human user’s dialog preference with a Elo rating system, while we focus on automatic and reproducible evaluation over popular benchmarks; 7) Chain-of-Thought Hub (Fu et al., 2023) focuses on evaluating the reasoning capability of LLMs with CoT prompting, while we support both CoT and answer-only prompting settings and evaluate various aspects of LLM capabilities.
@@ 2.2 Benchmarks for Evaluation
Safety. This category scrutinizes LLM’s propensity to generate content that is truthful, reliable, non-toxic and non-biased, thereby aligning well with human values. To this end, we currently have two sub-categories (and plan to support more benchmarks in the future): 1) Truthfulness, we employ TruthfulQA [^8] For TruthfulQA, we evaluate in the multiple-choice setting. (Lin et al., 2022), a benchmark designed to evaluate LLM’s factuality; 2) Toxicity, we adopt RealToxicityPrompts (Gehman et al., 2020) to quantify the risk of generating toxic output.
Note that the categories above are based on our own interpretation of LLM capabilities, which is by no means the exclusive approach to systematically evaluating LLMs. Additionally, some benchmarks may necessitate a range of capabilities. For instance, both “Knowledge” and “Reasoning” could influence the performance on MMLU. For the sake of simplicity, we just assign each benchmark to a primary capability category for evaluation. Due to the limited space, refer to Appendix B for details of dataset splits and source of prompts.
Apart from the capability categories listed above, we also plan to support more LLM capabilities in the future, such as long-context understanding, multi-turn conversation, open-domain generation, LLM-based autonomous agent, etc., some of which may require subjective assessment or using a powerful LLM to automatically evaluate LLMs such as Li et al. (2023).
2.3 Details of Black-box Evaluation
3 Experiments
Table 1: Main evaluation results of GPT-Fathom. Note that GPT-Fathom supports various settings for evaluation. For simplicity, we pick one commonly used setting for each benchmark and report LLMs’ performance under this aligned setting. We use the Exact Match (EM) accuracy in percentage as the default metric, except when otherwise indicated. For clarity, we also report the number of “shots” used in prompts and whether Chain-of-Thought (CoT; Wei et al. 2022) prompting is used. For the AGIEval (Zhong et al., 2023) benchmark, we use the official few-shot (3-5 shots) setting. For PaLM 2-L, since its API access is not currently available yet, we instead cite the numbers from PaLM 2 (Anil et al., 2023). Numbers that are not from our own experiments are shown in brackets. Numbers with ${}^{\star}$ are obtained from optimized prompts, which is discussed in Section 6.
@@ 3 Experiments
TriviaQA 1-shot 73.4 ${}^{\star}$ (86.1) 61.5 51.6 68.0 82.6 78.6 82.5 83.2 84.9 87.2 84.0 92.3 92.1 92.6
Multi-subject Test MMLU 5-shot ${}^{\star}$ ${}^{\star}$ (78.3) 34.3 39.9 46.7 69.1 62.1 63.7 66.6 67.4 69.6 61.9 83.7 81.3 78.3
@@
3.1 Overall Performance
From the overall performance of OpenAI’s models, we observe a remarkable leap from GPT-3 to GPT-4 across all facets of capabilities, with the GPT-3.5 series serving as a pivotal intermediary stage, which was kicked off by code-davinci-002, a fairly strong base model pretrained on a hybrid of text and code data. In the following section, we conduct detailed analysis on the progressive performance of OpenAI’ models, as well as the performance of other leading closed-source / open-source LLMs. Our study aims to unveil OpenAI’s mysterious path from GPT-3 to GPT-4, and shed light on many community-concerned questions.
@@ 3.2 Analysis and Insights
Caveats The analysis below is based on our benchmarking results and publicly available information of the evaluated LLMs. As claimed in OpenAI’s model index3, their models generally used the best available datasets at the time of training. As a result, our analysis on OpenAI’s models may not serve as a rigorous ablation study.
@@
Closed-source vs. open-source LLMs. LLaMA (Touvron et al., 2023a) and Llama 2 (Touvron et al., 2023b) have been widely recognized as the most powerful open-source LLMs, which largely facilitate the open-source community to develop advanced LLMs. Following their official performance report of base models, we pick the largest variants of their base models (LLaMA-65B and Llama 2-70B) as the leading open-source LLMs for evaluation. Compared to LLaMA, Llama 2 is trained on 40% more pretraining data with doubled context length (Touvron et al., 2023b). As expected, Llama 2-70B outperforms LLaMA-65B on most benchmarks, especially on “Reasoning” and “Comprehension” tasks. The radar chart in Figure 1(c) highlights the capability distribution of Llama 2-70B, which achieves similar performance on “Safety” against gpt-3.5-turbo-0613, but still clearly underperforms on the other dimensions, especially “Math”, “Coding” and “Multilingual”. We strongly encourage the open-source community to improve these capabilities of open-source LLMs.
@@
Impacts of pretraining with code data. Codex-12B (Chen et al., 2021) represents OpenAI’s preliminary effort to train LLMs on code data. Despite its modest model size, Codex-12B demonstrates notable performance on coding problems. Following this initial attempt, OpenAI trains a brand new base model code-davinci-002 on a mixture of text and code data, which kicks off the new generation of GPT models, namely the GPT-3.5 Series. As reported in Table 1, the performance of code-davinci-002 surges on all capability categories, compared to the GPT-3 Series, which is also visualized in Figure 1(a) and 1(b). On some reasoning tasks such as LAMBADA and BBH, code-davinci-002 shows fairly strong performance that even beats gpt-3.5-turbo-0301 and gpt-3.5-turbo-0613. This suggests that incorporating code data into LLM pretraining could universally elevate its potential, particularly in the capability of reasoning.
Table 3: Breakdown of coding performance with temperature $T=0.8$ and $\text{top}_{p}=1.0$.
@@
Table 4: Ablation study on number of “shots”.
Table 5: Ablation study on CoT prompting.
@@
Sampling variance. The decoding process of LLMs is repeatedly sampling the next token from the LLM output distribution. Various hyperparameters, including the temperature $T$ and the nucleus sampling (Holtzman et al., 2020) parameter $\text{top}_{p}$, can be adjusted to modify the sampling behavior. In our evaluations, we set $\text{top}_{p}=1.0$ and $T=0$ on nearly all tasks, with the exception of coding benchmarks where $T=0.8$. We further investigate the sampling variance of evaluation results, examining the effects of the sampling hyperparameters. Due to the limited space, in Appendix D, we report the mean and stand deviation of benchmark scores over 3 runs with different settings of $T$ and $\text{top}_{p}$. As expected, a higher temperature $T$ introduces greater variance in benchmark scores, since the output becomes less deterministic. Notably, LLMs (especially base models) tend to underperform with a higher temperature $T$. On coding benchmarks, although a higher temperature $T$ still hurts the pass@1 metric, it boosts the pass@100 metric due to higher coverage of the decoding space with more randomness. As for $\text{top}_{p}$, our results indicate that it has marginal influence on the performance of fine-tuned LLMs. Similarly, a notable exception is observed on coding benchmarks, where a higher $\text{top}_{p}$ diminishes the pass@1 metric but largely enhances the pass@100 metric.
4 Conclusions and Future Work
In this work, we present GPT-Fathom, an open-source and reproducible evaluation suite that comprehensively measures the multi-dimensional capabilities of LLMs under aligned settings. Our retrospective study on OpenAI’s models helps the community better understand the evolutionary path from GPT-3 to GPT-4, and sheds light on many questions that the community is eager to explore, such as the gap between leading closed-source / open-source LLMs, the benefits of pretraining with code data, the impacts of SFT and RLHF, etc. Moreover, we identify novel challenges of advanced LLMs, such as prompt sensitivity and the seesaw phenomenon of LLM capabilities.
In the future, we plan to further extend GPT-Fathom by 1) adding additional evaluation benchmarks under existing capability categories; 2) supporting more capability aspects, such as long-context understanding, multi-turn conversation, open-domain generation, LLM agent and even multi-modal capability; 3) evaluating more leading LLMs, including both open-source and closed-source models.
Acknowledgments. The authors would like to thank Yao Fu for the suggestions on benchmark selection. We also thank Ke Shen, Kai Hua, Yang Liu and Guang Yang for technical discussions. We gratefully acknowledge the funding support and feedback from Liang Xiang. This work was made possible by all the benchmarks used for evaluation. We appreciate the creators of these benchmarks.
@@ C.1 Sampling Hyperparameters
For coding evaluations, we sample 100 responses per question with temperature $T=0.8$. For all the other evaluations, we use $T=0$. The default $\text{top}_{p}=1.0$ is applied across all of our evaluations.
@@ C.2 Evaluation Prompts
AGIEval - Chinese MC 回答下列选择题
AGIEval - Chinese IMC (Indefinite MC) 回答下列多选题
AGIEval - Chinese Cloze 回答下列填空题
@@
C-Eval 以下是中国关于{task name}考试的单项选择题,请选出其中的正确答案。
@@ C.3 Answer Parsing and Metric Computation
We check if the model’s output aligns with any of the provided candidate answers.
@@
The F1 score is employed to measure performance.
@@
Appendix D Details of Experiments
In Table 9 and 10, we report the mean and stand deviation of benchmark scores over 3 runs, with different settings of $T$ and $\text{top}_{p}$.
Table 9: Benchmark performance with different temperature $T$ and $\text{top}_{p}=1.0$. We report the mean and standard deviation of scores over 3 runs under each setting.
@@ Appendix D Details of Experiments
Table 10: Benchmark performance with different temperature $T$ and $\text{top}_{p}$. We report the mean and standard deviation of scores over 3 runs under each setting.
Benchmark Setting      $\text{top}_{p}$ code-davinci-002 text-davinci-003 gpt-3.5-turbo-0301
@@ Appendix E Complete Results of LLaMA / Llama 2 Family
Table 12: Comparison of our evaluation results and GPT-4 officially reported scores. The official score of MATH is obtained from Bubeck et al. (2023), which is marked with ${}^{\star}$.

### v6

@@ 1 Introduction
  • Inconsistent settings: The evaluation settings, such as the number of in-context example “shots”, whether Chain-of-Thought (CoT; Wei et al. 2022) prompting is used, methods of answer parsing and metric computation, etc., often differ across the existing LLM works. Moreover, most of the released LLMs do not disclose their prompts used for evaluation, making it difficult to reproduce the reported scores. Different settings and prompts may lead to very different evaluation results, which may easily skew the observations. Yet, many existing LLM leaderboards reference scores from other papers without consistent settings and prompts, which may inadvertently encourage cherry-picking favored settings and prompts for better results. To achieve reliable conclusions, it is crucial to make apples-to-apples comparisons with consistent settings and prompts.
@@
2 Related Work
Benchmarks constantly play a pivotal role in steering the evolution of AI and, of course, directing the advancement of LLMs as well. There are many great existing LLM evaluation suites. By comparing GPT-Fathom with previous works, we summarize the major difference as follows: 1) HELM (Liang et al., 2023) primarily uses answer-only prompting (without CoT) and has not included the latest leading models such as GPT-4 (as of the time of writing); 2) Open LLM Leaderboard (Beeching et al., 2023) focuses on open-source LLMs, while we jointly consider leading closed-source and open-source LLMs; 3) OpenCompass (Contributors, 2023) evaluates latest open-source and closed-source LLMs (all released after 2023/03), while we cover both leading LLMs and OpenAI’s earlier models to decipher the evolutionary path from GPT-3 to GPT-4; 4) InstructEval (Chia et al., 2023) is designed for evaluating instruction-tuned LLMs, while we evaluate both base and SFT / RLHF models; 5) AlpacaEval (Li et al., 2023) evaluates on simple instruction-following tasks as a quick and cheap proxy of human evaluation, while we provide systematic evaluation of various aspects of LLM capabilities; 6) Chatbot Arena (Zheng et al., 2023) evaluates human user’s dialog preference with a Elo rating system, while we focus on automatic and reproducible evaluation over popular benchmarks; 7) Chain-of-Thought Hub (Fu et al., 2023) focuses on evaluating the reasoning capability of LLMs with CoT prompting, while we support both CoT and answer-only prompting settings and evaluate various aspects of LLM capabilities. We discuss more related work in Appendix G.
@@ 3.2 Benchmarks for Evaluation
Safety. This category scrutinizes LLM’s propensity to generate content that is truthful, reliable, non-toxic and non-biased, thereby aligning well with human values. To this end, we currently have two sub-categories: 1) Truthfulness, we employ TruthfulQA [^8] For TruthfulQA, we evaluate in the multiple-choice setting. (Lin et al., 2022), a benchmark designed to evaluate LLM’s factuality; 2) Toxicity, we adopt RealToxicityPrompts (Gehman et al., 2020) to quantify the risk of generating toxic output.
@@
TriviaQA 1-shot 73.4 74.0 (86.1) 61.5 51.6 68.0 82.6 78.6 82.5 83.2 84.9 87.2 84.0 92.3 92.1 92.6
Multi-subject Test MMLU 5-shot 60.1 67.8 (78.3) 34.3 39.9 46.7 69.1 62.1 63.7 66.6 67.4 69.6 61.9 83.7 81.3 78.3
@@
Table 1: Main evaluation results of GPT-Fathom. Note that GPT-Fathom supports various settings for evaluation. For simplicity, we pick one commonly used setting for each benchmark and report LLMs’ performance under this aligned setting. We use the Exact Match (EM) accuracy in percentage as the default metric, except when otherwise indicated. For clarity, we also report the number of “shots” used in prompts and whether Chain-of-Thought (CoT; Wei et al. 2022) prompting is used. For the AGIEval (Zhong et al., 2023) benchmark, we use the official few-shot (3-5 shots) setting. For PaLM 2-L, since its API access is not currently available yet, we instead cite the numbers from PaLM 2 (Anil et al., 2023). Numbers that are not from our own experiments are shown in brackets. Numbers with are obtained from optimized prompts, which is discussed in Section 6. Best and second best scores are highlighted in bold.
@@
3.3 Details of Black-box Evaluation
4 Experiments
4.1 Overall Performance
From the overall performance of OpenAI’s models, we observe a remarkable leap from GPT-3 to GPT-4 across all facets of capabilities, with the GPT-3.5 series serving as a pivotal intermediary stage, which was kicked off by code-davinci-002, a fairly strong base model pretrained on text and code data. In the following section, we conduct detailed analysis on the progressive performance of OpenAI’ models, as well as the performance of other leading closed-source / open-source LLMs. Our study aims to unveil OpenAI’s mysterious path from GPT-3 to GPT-4, and shed light on community-concerned questions.
@@ 4.2 Analysis and Insights
Closed-source vs. open-source LLMs. Recently, LLaMA (Touvron et al., 2023a) and Llama 2 (Touvron et al., 2023b) have been widely recognized as leading open-source LLMs, which largely facilitate the open-source community to develop advanced LLMs. Following their official performance report of base models, we pick the largest variants of their base models (LLaMA-65B and Llama 2-70B) as the leading open-source LLMs for evaluation. Compared to LLaMA, Llama 2 is trained on 40% more pretraining data with doubled context length (Touvron et al., 2023b). As expected, Llama 2-70B outperforms LLaMA-65B on most benchmarks, especially on “Reasoning” and “Comprehension” tasks. The radar chart in Figure 2(c) highlights the capability distribution of Llama 2-70B, which achieves similar performance on “Safety” against gpt-3.5-turbo-0613, but still clearly underperforms on the other dimensions, especially “Math”, “Coding” and “Multilingual”. We strongly encourage the open-source community to improve these capabilities of open-source LLMs.
@@
Impacts of pretraining with code data. Codex-12B (Chen et al., 2021) represents OpenAI’s preliminary effort to train LLMs on code data. Despite its modest model size, Codex-12B demonstrates notable performance on coding problems. Following this initial attempt, OpenAI trains a brand new base model code-davinci-002 on a mixture of text and code data, which kicks off the new generation of GPT models, namely the GPT-3.5 Series. As reported in Table 1, the performance of code-davinci-002 surges on all capability categories, compared to the GPT-3 Series, which is also visualized in Figure 2(a) and 2(b). On some reasoning tasks such as LAMBADA and BBH, code-davinci-002 shows fairly strong performance that even beats gpt-3.5-turbo-0301 and gpt-3.5-turbo-0613. This suggests that incorporating code data into LLM pretraining could universally elevate its potential, particularly in the capability of reasoning.
@@
TriviaQA 1-shot <q1>\$\backslash n$Answer: <a1>\$\backslash n$<q>\$\backslash n$Answer: 75.4 74.0 82.9 77.6 81.6 77.8 92.0
Q: <q1>\$\backslash n$A: <a1>\$\backslash n$Q: <q>\$\backslash n$A: 73.4 55.5 82.6 78.6 82.5 83.2 92.3
MMLU 5-shot <q1>\$\backslash n$Answer: <a1>\$\backslash n$ldots <q5>\$\backslash n$Answer: <a5>\$\backslash n$<q>\$\backslash n$Answer: 60.1 67.8 68.3 64.5 65.3 67.7 82.0
Q: <q1>\$\backslash n$A: <a1>\$\backslash n$ldots Q: <q5>\$\backslash n$A: <a5>\$\backslash n$Q: <q>\$\backslash n$A: 55.7 64.8 68.3 63.5 65.4 66.6 83.7
@@
Sampling variance. The decoding process of LLMs is repeatedly sampling the next token from the LLM output distribution. Various hyperparameters, including the temperature $T$ and the nucleus sampling (Holtzman et al., 2020) parameter topp, can be adjusted to modify the sampling behavior. In our evaluations, we set $top_{p}=1.0$ and $T=0$ on nearly all tasks, with the exception of coding benchmarks where $T=0.8$. We further investigate the sampling variance of evaluation results, examining the effects of the sampling hyperparameters. Due to the limited space, in Appendix D, we report the mean and stand deviation of benchmark scores over 3 runs with different settings of $T$ and topp. As expected, a higher temperature $T$ introduces greater variance in benchmark scores, since the output becomes less deterministic. Notably, LLMs (especially base models) tend to underperform with a higher temperature $T$. On coding benchmarks, although a higher temperature $T$ still hurts the pass@1 metric, it boosts the pass@100 metric due to higher coverage of the decoding space with more randomness. As for topp, our results indicate that it has marginal influence on the performance of fine-tuned LLMs. Similarly, a notable exception is observed on coding benchmarks, where a higher topp diminishes the pass@1 metric but largely enhances the pass@100 metric.
5 Conclusions
We present GPT-Fathom, an open-source and reproducible evaluation suite that comprehensively measures the multi-dimensional capabilities of LLMs under aligned settings. Our retrospective study on OpenAI’s models helps the community better understand the evolutionary path from GPT-3 to GPT-4, and sheds light on many community-concerned questions. For example, our study reveals that SFT and RLHF yields more benefit for weaker models, while for stronger base models, it offers diminishing returns or incurs an alignment tax. Moreover, we identify novel challenges of advanced LLMs, such as prompt sensitivity and the seesaw phenomenon of LLM capabilities.
Acknowledgments
The authors would like to thank Yao Fu for the suggestions on benchmark selection. We also thank Ke Shen, Kai Hua, Yang Liu and Guang Yang for technical discussions. We gratefully acknowledge the funding support and feedback from Liang Xiang. This work was made possible by all the benchmarks used for evaluation. We appreciate the creators of these benchmarks.
Limitations
While this work brings forth novel insights on LLM evaluation, it presents certain limitation. Primarily, we rely on regular expression matching to extract answers from LLMs’ responses. While this method proves to be effective in most instances, in some corner cases, it might overlook the actual answers provided by the models, particularly when the evaluated LLM has a limited capacity to adhere to the instructions. A possible solution to this issue involves the utilization of more sophisticated LLMs, such as GPT-4, for answer extraction. However, this approach may also incur significant costs. We hope that continued advancements in LLMs evaluation will improve the answering parsing techniques.
@@ C.1 Sampling Hyperparameters
[…]
@@ C.2 Evaluation Prompts
AGIEval - Chinese MC {CJK}UTF8gbsn回答下列选择题
AGIEval - Chinese IMC (Indefinite MC) {CJK}UTF8gbsn回答下列多选题
AGIEval - Chinese Cloze {CJK}UTF8gbsn回答下列填空题
@@
C-Eval {CJK}UTF8gbsn以下是中国关于{task name}考试的单项选择题,请选出其中的正确答案。
@@ C.3 Answer Parsing and Metric Computation
The model’s performance is gauged using the F1 score, without any post-processing such as case normalization.
@@
Similarly, the F1 score is employed to measure performance.
@@
Appendix D Sampling Variance
In Table 9 and 10, we report the mean and stand deviation of benchmark scores over 3 runs, with different settings of $T$ and topp.
@@ Appendix D Sampling Variance
[…]
Benchmark Setting      topp code-davinci-002 text-davinci-003 gpt-3.5-turbo-0301
@@
Table 10: Benchmark performance with different temperature $T$ and topp. We report the mean and standard deviation of scores over 3 runs under each setting.
@@ Appendix E Complete Results of LLaMA / Llama 2 Family
Table 12: Comparison of our evaluation results and GPT-4 officially reported scores. The official score of MATH is obtained from Bubeck et al. (2023), which is marked with .
@@
Appendix G Related Work on Analysis
The impacts of pre-training on code data, along with the effects of Supervised Fine-Tuning (SFT) and Reinforcement Learning from Human Feedback (RLHF) on Large Language Models (LLMs), represent significant areas of research interest, with several concurrent studies exploring these themes. Notably, the work by Ma et al. (2023) and Yang et al. (2024) underscores the benefits of integrating code data into the pre-training phase for LLMs, demonstrating an improvement in coding and reasoning abilities without compromising performance on other tasks. Furthermore, the application of SFT and RLHF methods has been shown to reduce instances of hallucination Li et al. (2024), enhance output diversity and generalization Kirk et al. (2024), and improve safety measures Lin et al. (2023); Shen et al. (2024). However, these advancements come at the cost of an alignment tax, a challenge also recognized in the literature Askell et al. (2021); Liu et al. (2022). Our research primarily investigates these areas through a comprehensive evaluation framework, aligning with and potentially corroborating the findings of these studies.
Appendix H Impacts of In-context Samples
Benchmark Setting gpt-3.5-turbo-0613 gpt-4-0613 Llama 2-70B
ARC-c Sample Selection 1 81.6 84.6 79.6
Sample Selection 2 82..3 95.2 80.2
Sample Selection 3 81.7 95.0 82.2
DROP Sample Selection 1, F1 53.4 74.4 12.1
Sample Selection 2, F1 46.4 76.2 15.4
Table 14: Impacts of the selection of in-context samples.
To investigate the impact of in-context sample selection on performance, we carried out experiments on ARC-c and DROP, utilizing different sets of in-context learning samples. The findings, presented in Table 14, indicate that the choice of in-context samples indeed affects the model’s performance. While the influence on multiple-choice questions is minimal, for free-form questions, the impact of few-shot samples becomes more pronounced. This underscores the significance of disclosing the few-shot samples employed during evaluation.
Appendix I Language-wise Evaluation Results
Language-wise evaluation results for multilingual tasks are summarized in Table 15 and Table 16.
Language davinci davinci- instruct-beta (InstructGPT) text davinci -001 code- davinci -002 text- davinci- 002 text- davinci- 003 gpt-3.5- turbo- 0314 gpt-3.5- turbo- 0613 gpt-4- 0314 gpt-4- 0613
LLaMA-
65B
Llama 2-
70B
Bengali 1.60 1.20 3.20 9.60 6.00 15.60 35.20 38.80 66.40 34.40 4.00 4.00
German 3.60 12.00 14.40 6.40 32.00 52.80 71.20 68.00 87.20 87.20 1.60 6.80
Spanish 0.80 10.40 12.40 4.00 38.80 54.40 74.00 70.40 88.80 90.00 5.60 6.00
French 3.60 12.00 11.20 4.80 42.00 54.40 67.60 69.20 82.40 82.80 6.40 5.60
Japanese 4.80 2.80 10.00 7.60 32.80 39.20 60.40 46.00 82.80 83.20 2.80 5.20
Russian 2.80 3.20 8.40 23.60 26.40 40.40 70.40 67.20 87.20 86.80 2.80 3.60
Swahili 2.40 3.60 3.20 8.40 13.20 26.00 56.80 56.00 84.40 84.40 2.40 2.80
Telugu 0.80 0.40 1.60 1.60 0.80 1.20 2.00 16.40 10.40 4.80 0.80 1.20
Thai 0.40 1.20 3.20 4.80 6.00 6.00 35.60 45.20 53.60 48.40 3.60 3.20
Chinese 3.60 4.40 6.00 8.00 31.20 47.20 61.60 59.60 82.40 84.80 6.40 1.60
Table 15: Language-wise evaluation results on MGSM.
Language davinci davinci- instruct-beta (InstructGPT) text davinci -001 code- davinci -002 text- davinci- 002 text- davinci- 003 gpt-3.5- turbo- 0314 gpt-3.5- turbo- 0613 gpt-4- 0314 gpt-4- 0613
LLaMA-
65B
Llama 2-
70B
arabic 1.80 1.80 0.04 0.08 0.06 0.08 0.22 23.80 28.10 28.70 7.10 11.90
bengali 0.00 2.70 0.02 0.04 0.04 0.04 0.15 20.40 23.00 23.00 0.90 7.10
english 19.80 12.30 0.28 0.26 0.28 0.38 0.33 36.40 43.40 43.40 23.90 20.00
finnish 14.80 5.80 0.18 0.27 0.22 0.27 0.35 36.70 39.10 39.00 20.80 34.30
indonesian 10.30 3.00 0.18 0.19 0.15 0.24 0.25 28.30 32.60 32.70 17.50 22.80
japanese 5.50 6.40 0.10 0.18 0.16 0.19 0.31 34.70 43.10 42.60 15.20 33.40
korean 2.20 0.70 0.06 0.09 0.10 0.19 0.31 27.20 41.70 41.70 11.20 33.70
russian 3.60 3.70 0.05 0.10 0.06 0.09 0.11 17.60 23.20 23.40 14.70 16.30
swahili 2.80 1.80 0.10 0.31 0.28 0.28 0.26 40.30 49.70 48.70 20.40 23.20
telugu 0.00 0.00 0.00 0.00 0.00 0.01 0.02 2.70 5.70 5.70 0.00 0.90
thai 2.00 2.50 0.02 0.05 0.03 0.04 0.04 7.60 14.40 14.60 1.80 3.20
Table 16: Language-wise evaluation results on TyDi QA.